How would you rearrange the folloowing equation : 3x-2y=15 ? I dont need to solve it, just how to rearrange it?
Answer by drk(1908)

You could rearrange so that x is negative and y is positive as
-3x + 2y = 15
---
You could rearrange so that you solve for x as
x = (2/3)y + 5
---
You could rearrange so that you solve for y as
y = (3/2)y -15/2
however, you said not to solve it.
So, I am not sure which direction to go.